### v1

@@
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] BRAIn Lab    [Uncaptioned image] WeightLoRA: Keep Only Necessary Adapters Andrey Veprikov1, 2, 3, Vladimir Solodkin1, 2, 3, Alexander Zyl1, Andrey Savchenko3, Aleksandr Beznosikov1, 2, 4 1Moscow Institute of Physics and Technology 2Institute for System Programming, RAS 3Sber AI Lab 4Innopolis University The widespread utilization of language models in modern applications is inconceivable without Parameter-Efficient Fine-Tuning techniques, such as low-rank adaptation (LoRA), which adds trainable adapters to selected layers. Although LoRA may obtain accurate solutions, it requires significant memory to train large models and intuition on which layers to add adapters. In this paper, we propose a novel method, WeightLoRA, which overcomes this issue by adaptive selection of the most critical LoRA heads throughout the optimization process. As a result, we can significantly reduce the number of trainable parameters while maintaining the capability to obtain consistent or even superior metric values. We conduct experiments for a series of competitive benchmarks and DeBERTa, BART, and Llama models, comparing our method with different adaptive approaches. The experimental results demonstrate the efficacy of WeightLoRA and the superior performance of WeightLoRA+ in almost all cases.
@@ 1   Introduction
The emergence of Large Language Models (LLMs) has transformed numerous domains, from natural language processing to code generation [Dubey et al., 2024; He et al., 2020]. Despite the undebatable performance, their fine-tuning remains computationally intensive and memory-demanding. Parameter-Efficient Fine-Tuning (PEFT) [Ding et al., 2023; Han et al., 2024], such as Low-Rank Adaptation (LoRA[Hu et al., 2021], has significantly alleviated these challenges by reducing the number of trainable parameters while maintaining competitive accuracy.
@@ 2   Related Work
DoRA (Weight-Decomposed LoRA[Liu et al., 2024a] decomposes the LLM’s weight into magnitude and direction parts. While it achieves competitive performance, it introduces an overhead in training time, which can be undesirable. Another modification, NoRA [Lin et al., 2024a] uses singular value decomposition (SVD) and a special dual structure, where the inner layer introduces fine-grained adjustments tailored to different tasks, and the outer layer provides a low-rank foundation by keeping its weights fixed. MiLoRA [Wang et al., 2024] performs SVD and trains only the minor singular components.
@@
Allocating LoRA (ALoRA[Liu et al., 2024b] utilizes the diagonal matrix of gate units for each rank not greater than the initial LoRA rank. During training, the rank gates are set to zero for ranks with negative contributions, and the pruned rank budget is added to other weights. As a result, it can increase the accuracy of fine-tuning for a fixed memory budget.
@@ 3.1 Motivating Example
(1)$\Leftrightarrow&\leavevmode\nobreak\ \alpha_{1}\min_{A^{1}B^{1}}\left\{\langle\nabla_{W^{1}}f(\mathcal{W}),A^{1}B^{1}\rangle\right\}+\ldots\\ &+\alpha_{n}\min_{A^{n}B^{n}}\left\{\langle\nabla_{W^{n}}f(\mathcal{W}),A^{n}B^{n}\rangle\right\},\end{split}$ (2)
In the general case, given this setting, all minima are distinct and have different magnitudes, illustrating potential differences in the tuning process. Consequently, different times may be required to optimize different $A^{i}B^{i}$ during training. Furthermore, it is possible that $A^{i}B^{i}$ does not need to be tuned at all, e.g., when the condition $\nabla_{W^{i}}f(W^{1},\ldots,W^{n})=0$ is met, which means that we do not need any fine-tuning for the layer $i$.
@@ 3.3 WeightLoRA
[…]
@@
It is important to note that simply adding the weights $\omega_{i}$ to the matrices $A^{i}B^{i}$ does not make much sense, given that their parameters are already trainable. That is why incorporating the $l_{0}$-norm is a fundamental component of our method. Due to the same trainability, the characteristic values of the weights $w_{i}$ are expected to be either 00 or $1$, aligning with the observations made in numerical experiments (see Section 4).
@@ 3.4 WeightLoRA+
  • Extend existing matrices as it was done in the original paper [Hu et al., 2021], i.e., $A$ is expanded via a random Gaussian initialization and $B$ is expanded with zeros:
@@ 4.2 Question Answering
Table 4: Results of fine-tuning DeBERTaV3-base on SQuAD v1.1 and SQuAD v2.0 datasets. We report F1-scores averaged by $5$ runs. The best results are shown in bold, the second-best are underlined. All experiment details are provided in Appendix B.2.
@@ 4.3 Natural Language Generation
Table 5: Results of fine-tuning BART-large on XSum and CNN/DailyMail datasets. We report rogue1-scores averaged by $5$ runs. The best results are shown in bold, the second-best are underlined. All experiment details are provided in Appendix B.2.
@@ Appendix A LoRA Notation
Let us briefly describe the original LoRA [Hu et al., 2021] framework. Given a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA aims to learn an additive update $\Delta W\in\mathbb{R}^{d\times k}$, which is constrained to a low-rank decomposition of the form: $\Delta W=AB$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$. It is assumed that the rank $r$ is relatively small, $r\ll\min(d,k)$, to reduce the number of trainable parameters. This low-rank decomposition enables task-specific information to be encoded in the matrices $A$ and $B$ without requiring fine-tuning of the entire $W$. Formally, given an input $x$, the forward of the layer $h$ is as follows:
@@
In this paper, we also define the sets of all adapters $A$ and $B$ (for all layers) as $\mathcal{A}$ and $\mathcal{B}$, respectively.
@@ Appendix B Experimental Details
We use PyTorch [Paszke et al., 2019] to implement our algorithms and the publicly available implementations of the baselines. Our code is based on the publicly available Huggingface Transformers3 [Wolf et al., 2020] library. All the experiments are conducted on NVIDIA V100 16GB and NVIDIA A100 80GB GPU. All the results are averaged across $5$ random seeds.
@@
[…]
LLama3 8B experiment. We tune the learning rate using the same grid as in the previous experiment. Warmup steps were tuned on the set $\{10,25,30,50\}$. The number of training steps and the batch size are adjusted according to the model size. In this experiment, we also use $2$ gradient accumulation steps instead of $6$, as was done in both GLUE experiments. The batch sizes and number of training steps for each task are provided in Table 8.

### v2

@@
BRAIn Lab[Uncaptioned image] WeightLoRA: Keep Only Necessary Adapters Andrey Veprikov1, Vladimir Solodkin1, Alexander Zyl1, Andrey Savchenko2, Aleksandr Beznosikov1, 3 1Moscow Center for Advanced Studies 2SB AI Lab 3Innopolis University The widespread utilization of language models in modern applications is inconceivable without Parameter-Efficient Fine-Tuning techniques, such as low-rank adaptation (LoRA), which adds trainable adapters to selected layers. Although LoRA may obtain accurate solutions, it requires significant memory to train large models and intuition on which layers to add adapters. In this paper, we propose a novel method, WeightLoRA, which overcomes this issue by adaptive selection of the most critical LoRA heads throughout the optimization process. As a result, we can significantly reduce the number of trainable parameters while maintaining the capability to obtain consistent or even superior metric values. We conduct experiments for a series of competitive benchmarks and DeBERTa, BART, and Llama models, comparing our method with different adaptive approaches. The experimental results demonstrate the efficacy of WeightLoRA and the superior performance of WeightLoRA+ in almost all cases.
@@ 1  Introduction
The emergence of Large Language Models (LLMs) has transformed numerous domains, from natural language processing to code generation [Dubey et al., 2024; He et al., 2020]. Despite the undebatable performance, their fine-tuning remains computationally intensive and memory-demanding. Parameter-Efficient Fine-Tuning (PEFT) [Ding et al., 2023; Han et al., 2024], such as Low-Rank Adaptation (LoRA) [Hu et al., 2021], has significantly alleviated these challenges by reducing the number of trainable parameters while maintaining competitive accuracy.
@@ 2  Related Work
DoRA (Weight-Decomposed LoRA) [Liu et al., 2024a] decomposes the LLM’s weight into magnitude and direction parts. While it achieves competitive performance, it introduces an overhead in training time, which can be undesirable. Another modification, NoRA [Lin et al., 2024a] uses singular value decomposition (SVD) and a special dual structure, where the inner layer introduces fine-grained adjustments tailored to different tasks, and the outer layer provides a low-rank foundation by keeping its weights fixed. MiLoRA [Wang et al., 2024] performs SVD and trains only the minor singular components.
@@
Allocating LoRA (ALoRA) [Liu et al., 2024b] utilizes the diagonal matrix of gate units for each rank not greater than the initial LoRA rank. During training, the rank gates are set to zero for ranks with negative contributions, and the pruned rank budget is added to other weights. As a result, it can increase the accuracy of fine-tuning for a fixed memory budget.
@@ 3.1 Motivating Example
[…]
In the general case, given this setting, all minima are distinct and have different magnitudes, illustrating potential differences in the tuning process. Consequently, different times may be required to optimize different $A^{i}B^{i}$ during training. Furthermore, it is possible that $A^{i}B^{i}$ does not need to be tuned at all, e.g., when the condition $\nabla_{W^{i}}f(W^{1},\ldots,W^{n})=0$ is met, which means that we do not need any fine-tuning for the layer ii.
@@ 3.3 WeightLoRA
where $\omega_{i}$ represents the importance weight of the $i$-th adapter. The weights are composed into a vector $\omega=(\omega_{1},\ldots,\omega_{n})^{T},$ which is then optimized throughout the fine-tuning process. To control the number of active adapters, we propose conditioning the $l_{0}$-‘‘norm’’ of $\omega$: $\|\omega\|_{0}\overset{\text{def}}{=}\sum\nolimits_{i=1}^{n}\mathbb{I}(\omega_{i}\neq 0)$.
@@
It is important to note that simply adding the weights $\omega_{i}$ to the matrices $A^{i}B^{i}$ does not make much sense, given that their parameters are already trainable. That is why incorporating the $l_{0}$-norm is a fundamental component of our method. Due to the same trainability, the characteristic values of the weights $w_{i}$ are expected to be either 0 or 11, aligning with the observations made in numerical experiments (see Section 4).
@@ 3.4 WeightLoRA+
  • Extend existing matrices as it was done in the original paper [Hu et al., 2021], i.e., AA is expanded via a random Gaussian initialization and $B$ is expanded with zeros:
@@ 4.2 Question Answering
Table 4: Results of fine-tuning DeBERTaV3-base on SQuAD v1.1 and SQuAD v2.0 datasets. We report F1-scores averaged by 55 runs. The best results are shown in bold, the second-best are underlined. All experiment details are provided in Appendix B.2.
@@ 4.3 Natural Language Generation
Table 5: Results of fine-tuning BART-large on XSum and CNN/DailyMail datasets. We report rogue1-scores averaged by 55 runs. The best results are shown in bold, the second-best are underlined. All experiment details are provided in Appendix B.2.
@@ Appendix A LoRA Notation
Let us briefly describe the original LoRA [Hu et al., 2021] framework. Given a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA aims to learn an additive update $\Delta W\in\mathbb{R}^{d\times k}$, which is constrained to a low-rank decomposition of the form: $\Delta W=AB$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$. It is assumed that the rank $r$ is relatively small, $r\ll\min(d,k)$, to reduce the number of trainable parameters. This low-rank decomposition enables task-specific information to be encoded in the matrices AA and $B$ without requiring fine-tuning of the entire $W$. Formally, given an input $x$, the forward of the layer $h$ is as follows:
@@
In this paper, we also define the sets of all adapters AA and $B$ (for all layers) as $\mathcal{A}$ and $\mathcal{B}$, respectively.
@@ Appendix B Experimental Details
We use PyTorch [Paszke et al., 2019] to implement our algorithms and the publicly available implementations of the baselines. Our code is based on the publicly available Huggingface Transformers3 [Wolf et al., 2020] library. All the experiments are conducted on NVIDIA V100 16GB and NVIDIA A100 80GB GPU. All the results are averaged across 55 random seeds.
@@
Dynamic rank comparison. We select the learning rate that achieves the best performance from the grid $\{5\cdot 10^{-5},\ 5\cdot 10^{-4},\ 5\cdot 10^{-3}\}$. We apply LoRA-like adapters only to the self-attention layers, while the other attention layers remain frozen. The total number of adapters for DeBERTaV3-base is $n=36$. For all GLUE tasks, we use one training epoch, a batch size of $32$, and 66 gradient accumulation steps. The scheduler uses a linear schedule with $100$ warm-up steps.
@@
LLama3 8B experiment. We tune the learning rate using the same grid as in the previous experiment. Warmup steps were tuned on the set $\{10,25,30,50\}$. The number of training steps and the batch size are adjusted according to the model size. In this experiment, we also use 22 gradient accumulation steps instead of 66, as was done in both GLUE experiments. The batch sizes and number of training steps for each task are provided in Table 8.